OpenStudy (anonymous):

PLEASE HELP MEDAL AND FAN I NEED 3 POINTS FOR : -X+2Y=11

Mehek (mehek14):

question?

OpenStudy (texaschic101):

pick a number, any number...plug it in for x, then solve for y

OpenStudy (anonymous):

^^^^^^^ not sure how to do so

OpenStudy (anonymous):

Find 3 points that solve the equation then plot on the graph @Mehek14

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

lets let x be 2 now sub 2 in for x in the equation -x + 2y = 11 -(2) + 2y = 11 -- add 2 to both sides 2y = 11 + 2 2y = 13 y = 13/2 or 6 1/2 so one point is : (2, 13/2)...because all points are set up as (x,y)

OpenStudy (texaschic101):

understand ?

OpenStudy (texaschic101):

for another point, you can let x be 3, and sub it into the equation to find y.

OpenStudy (anonymous):

still dont understand @texaschic101